Cellphone sales person wishes to find the 99% confidence interval estimate for the true number of customers who upgrade to a new plan each year. How large should their sample be if they wish to be within 5% of the true proportion?
Answer #1

Solution :

Given that,

= 0.5 ( assume 0.5)

1 - = 1 - 0.5 = 0.5

margin of error = E = 5% = 0.05

At 99% confidence level the z is,

= 1 - 99%

= 1 - 0.99 = 0.01

/2 = 0.005

Z/2 = 2.58    ( Using z table )

Sample size = n = (Z/2 / E)2 * * (1 - )

= (2.58 / 0.05)2 * 0.5 * 0.5

=665.64

Sample size =666